Write an equation for a sine wave with amplitude 0.120 m, wavelength 0.293 m, and wave speed 6.41 m/s traveling in the -x-direction. y(x,t) = (a m)sin[( b s-1)t + (c m-1)x] where a =

b =

c =
Answer #1

y(x,t) = (a m)sin[( b s-1)t + (c m-1)x]

In the standard form

a = amplitude

So,

a = 0.120 m

b = angular frequency

Now we know that

wave speed = frequency*wavelength

f = v/lambda = 6.41/0.293 = 21.88 Hz

Now,

angular frequency = 2*pi*f = 2*pi*21.88

b = 137.48 s^-1

c = wave number

wave number = 2*pi/lambda

= 2*pi/0.293 = 21.44 m^-1

c = 21.44 m^-1